OpenStudy (anonymous):

Rewrite in simplest rational exponent form √x • 4√x. Please show each step of your process.

OpenStudy (anonymous):

I'd appreciate any help:)

OpenStudy (anonymous):

Times them together...\( \sqrt{x} \times 4\sqrt{x} \) \( \sqrt{x} \times \sqrt{x} \) =\( \sqrt{x^2} \) = \( x \) then times x by 4, which gives you 4x

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok thank-you. Can i ask another?